# Continuity of extensions of Lipschitz maps and of monotone maps

### v1

@@ 4.2. Consequence of existence of isometric embedding
The strategy of the proof that iii) of Theorem 1.2 implies i) of that theorem that we adapt is the following. Let $v\colon X\to Y$. Let $x_{1},\dotsc,x_{m+1}\in X$. We shall construct an isometric embedding $u\colon\{x_{1},\dotsc,x_{m+1}\}\to Y$ in such a way that:
@@
    for any $i=1,2,\dotsc,m+1$, $u(x_{i})$ is roughly at a distance $\delta$ to the corresponding vertex $v(x_{i})$, $i=1,2,\dotsc,m$,
@@
    the simplex is shifted in direction of $-(v(x_{m+1})-\sum_{i=1}^{m}t_{i}v(x_{i}))$ with respect to $\mathrm{Conv}\{v(x_{1}),\dotsc,v(x_{m+1})\}$,
then i) of Theorem 1.2 is valid for $v$, provided that $\delta$ is sufficiently large. This conclusion is the matter of the following lemma.
@@
(4.1) $8(\mathrm{diam}A)^{2}+5C+\Bigg{(}\frac{10\,(\mathrm{diam}A)^{2}+3C}{\Big{\lVert}v(x_{m+1})-\sum_{i=1}^{m}t_{i}v(x_{i})\Big{\rVert}-\Big{\lVert}x_{m+1}-\sum_{i=1}^{m}t_{i}x_{i}\Big{\rVert}}\Bigg{)}^{2}.$
@@
and such that for some non-negative numbers $s_{1},\dotsc,s_{m}$ that sum up to one, some non-negative $\lambda$ there is
@@
$\bigg{\lvert}\Big{\lVert}\sum_{i=1}^{k}t_{i}(u(x_{i})-v(x_{i}))\Big{\rVert}^{2}-\delta\bigg{\lvert}\leq 8\big{(}\mathrm{diam}A\big{)}^{2}+5C.$
@@
$\Big{\lvert}2\Big{\langle}u(x_{i})-v(x_{i})+\big{(}u(x_{j})-v(x_{j})\big{)},u(x_{j})-v(x_{j})\Big{\rangle}-\lVert u(x_{i})-v(x_{i})-(u(x_{j})-v(x_{j}))\rVert^{2}\Big{\rvert}\leq 4C.$
Therefore for each $j=1,\dotsc,k$
@@
$\Big{\lvert}\Big{\lVert}\sum_{i=1}^{k}t_{i}(u(x_{i})-v(x_{i})-(u(x_{j})-v(x_{j})))\Big{\rVert}^{2}+\sum_{i=1}^{k}t_{i}\lVert u(x_{i})-v(x_{i})-(u(x_{j})-v(x_{j}))\rVert^{2}\Big{\rvert}+5C.$
@@
$\bigg{\lvert}\Big{\lVert}\sum_{i=1}^{k}t_{i}(u(x_{i})-v(x_{i}))\Big{\rVert}^{2}-\delta\bigg{\lvert}\leq 8\big{(}\mathrm{diam}A\big{)}^{2}+5C.$
@@
(4.3) $\bigg{\lvert}\Big{\lVert}\sum_{i=1}^{m}t_{i}(u(x_{i})-v(x_{i}))\Big{\rVert}^{2}-\delta\bigg{\lvert}\leq 8\big{(}\mathrm{diam}A\big{)}^{2}+5C.$
@@
$2\Big{\langle}\Delta_{u}-\Delta_{v},\sum_{i=1}^{m}t_{i}(u(x_{i})-v(x_{i}))\Big{\rangle}+\Big{\lVert}\sum_{i=1}^{m}t_{i}(u(x_{i})-v(x_{i}))\Big{\rVert}^{2}+\lVert\Delta_{u}-\Delta_{v}\rVert^{2}\leq\delta^{2}+C.$
@@
$=-2\lambda\Big{\langle}\Delta_{u}-\Delta_{v},\Delta_{v}\Big{\rangle}+2\Big{\langle}\Delta_{u}-\Delta_{v}\sum_{i,j=1}^{m}t_{i}s_{j}(u(x_{i})-v(x_{i})-u(x_{j})+v(x_{j}))\Big{\rangle}.$
Thus, thanks to $1$-Lipschitzness of $u$ and of $v$, and to (4.3), we get
$-2\lambda\Big{\langle}\Delta_{v},\Delta_{u}-\Delta_{v}\Big{\rangle}\leq 20(\mathrm{diam}A)^{2}+6C.$
Hence, by (4.5),
$2\lambda\lVert\Delta_{v}\rVert^{2}\leq 20(\mathrm{diam}A)^{2}+6C+2\lambda\lVert\Delta_{v}\rVert\lVert\Delta_{x}\rVert.$
@@
$\lambda\lVert\Delta_{v}\rVert\leq\frac{10(\mathrm{diam}A)^{2}+3C}{\lVert\Delta_{v}\rVert-\lVert\Delta_{x}\rVert}.$
Similarly as in (4.3) we show that, using (4.2)
$\lvert\lambda^{2}\lVert\Delta_{v}\rVert^{2}-\delta\rvert\leq 8(\mathrm{diam}A)^{2}+5C.$
@@
$\delta\leq\lambda^{2}\lVert\Delta_{v}\rVert^{2}+8(\mathrm{diam}A)^{2}+5C\leq 8(\mathrm{diam}A)^{2}+5C+\bigg{(}\frac{10\,(\mathrm{diam}A)^{2}+3C}{\lVert\Delta_{v}\rVert-\lVert\Delta_{x}\rVert}\bigg{)}^{2}.$
@@ 4.3. Existence of embedding and completion of the proof for $\mathrm{dim}Y\leq 3$
It will follow that (4.6) is also satisfied. Indeed, (4.7) implies that trivially that
@@
For any distinct numbers $i,j\in\{1,2,3\}$ we may find $u(x_{i}),u(x_{j})$ in the way described above. We infer by Lemma 4.4 that for any $j\in\{1,2,3\}$ and any non-negative real numbers $(r_{i})_{i\in\{1,2,3\}\setminus\{j\}}$ that sum up to one there is
$\Big{\lVert}v(x_{j})-\sum_{i\in\{1,2,3\}\setminus\{j\}}r_{i}v(x_{i})\Big{\rVert}\leq\Big{\lVert}x_{j}-\sum_{i\in\{1,2,3\}\setminus\{j\}}r_{i}x_{i}\Big{\rVert},$
@@
Next, we shall construct $u(x_{3})$, so that $u\colon\{x_{1},x_{2},x_{3}\}\to Y$ is an isometry and such that (4.7) holds true. Observe that the possible positions for $u(x_{3})$, for which $u\colon\{x_{1},x_{2},x_{3}\}\to Y$ is an isometry, lie on a circle contained in $Y$. The intersection of the considered circle with $W$ consists of two points $y_{+},y_{-}$. There are two possible positions of $v(x_{3})$ in $W$, given the positions of $v(x_{1}),v(x_{2})$, which we shall denote by $z_{+}$ and $z_{-}$. The situation is illustrated below.
@@
Using (4.12), we see that $\lVert v(x_{1})-v(x_{2})\rVert=\lVert x_{1}-x_{2}\rVert$. Thus, in (4.15), we get equalities, which also shows that $u(x_{3})$ that satisfies our requirements exists.
@@
Then $v$ satisfies i) of Theorem 1.2, cf. Remark 3.1. We shall specify $C>0$ and $\delta>0$ later. Suppose that we can find an isometry $u\colon A\to Y$ with required properties. Then
@@
$\lvert\langle u((1,1))-v((1,1)),w\rangle-\delta\rvert=\Big{\lvert}\langle u((1,1))-u((0,0)),w\rangle-\frac{1}{\sqrt{2}}\Big{\rvert}$
@@
Let us suppose that $X$ is a convex set and $\mathrm{dim}Y\geq 2$. Let $v\colon X\to\mathbb{R}^{m}$. Then $v$ is $1$-Lipschitz and affine if and only if for all finite $m\leq\mathrm{dim}Y$, all $x,x_{1},\dotsc,x_{m}\in X$ and all non-negative real numbers $t_{1},\dotsc,t_{m}$ that sum up to one there is
@@
In particular, if $K\subset X$ is a convex set and $\mathrm{dim}Y\geq 2$, then any $v$ satisfying (4.21) is affine on $K$.
@@
Let us suppose that $X$ is convex and that iii) holds true. By Lemma 4.4 and Lemma 4.8 we see that for $m\leq\min\{2,\mathrm{dim}Y\}$ and all $x,x_{1},\dotsc,x_{m}\in X$ and all non-negative real numbers $t_{1},\dotsc,t_{m}$ that sum up to one there is
@@ 5. Existence of extensions of monotone maps
We need to check that it is a $K$-function, see Definition 2.1. The condition of convexity and finitely lower semi-continuity is clearly satisfied. We need only to check whether the condition (2.1) holds, with $l\leq m$, as $\mathrm{dim}V\leq k-1\leq m-1$. Let $l\leq m$ and $(y_{i},x_{i})_{i=1}^{l}\in V\times X$, $x\in X$ and let $\lambda_{1},\dotsc,\lambda_{l}\geq 0$ be such that $\sum_{i=1}^{l}\lambda_{i}=1$. Then a short calculation readily implies that (2.1) for $\Phi$ is equivalent to
@@
By Theorem 2.2 we infer that for any points $x_{1},\dotsc,x_{k}\in A$, $k\leq m$, and any $x\in X$ the intersection (5.1) of closed, convex sets is non-empty. Indeed, condition that $\Phi(y-y_{i},x_{i},x)\leq 0$, for $i=1,\dotsc,k$ and $y_{i}=u(x_{i})-v(x_{i})$, is equivalent to that for $i=1,\dotsc,k$
$\langle y+v(x)-(y_{i}+v(x_{i})),x-x_{i}\rangle\geq 0.$
@@ 6. Further generalisations
One of the possible to obtain results concerns extensions of maps $u\colon A\to Y$ that satisfy

### v2

@@ 4.2. Consequence of existence of isometric embedding
The strategy of the proof that iii) of Theorem 1.2 implies i) of that theorem that we adopt is the following. Let $v\colon X\to Y$. Let $x_{1},\dotsc,x_{m}\in X$. We shall construct an isometric embedding $u\colon\{x_{1},\dotsc,x_{m}\}\to Y$ in such a way that:
@@
    for any $i=1,2,\dotsc,m$, $u(x_{i})$ is roughly at a distance $\delta$ to the corresponding vertex $v(x_{i})$,
@@
    the simplex $\mathrm{Conv}\{u(x_{1}),\dotsc,u(x_{m})\}$ is shifted in direction of
    $v(x_{m+1})-\sum_{i=1}^{m}t_{i}v(x_{i})$
    with respect to $\mathrm{Conv}\{v(x_{1}),\dotsc,v(x_{m})\}$.
Then we can infer that i) of Theorem 1.2 is valid for $v$, provided that we pick $\delta$ sufficiently large. This conclusion is the matter of the following lemma.
@@
(4.1) $8(\mathrm{diam}A)^{2}+3C+\Bigg{(}\frac{8\,(\mathrm{diam}A)^{2}+2C}{\Big{\lVert}v(x_{m+1})-\sum_{i=1}^{m}t_{i}v(x_{i})\Big{\rVert}-\Big{\lVert}x_{m+1}-\sum_{i=1}^{m}t_{i}x_{i}\Big{\rVert}}\Bigg{)}^{2}.$
@@
and such that for some non-negative numbers $s_{1},\dotsc,s_{m}$ that sum up to one, some non-negative $\lambda$ we have
@@
$\bigg{\lvert}\Big{\lVert}\sum_{i=1}^{k}t_{i}(u(x_{i})-v(x_{i}))\Big{\rVert}^{2}-\delta\bigg{\lvert}\leq 8\big{(}\mathrm{diam}A\big{)}^{2}+3C.$
@@
$\Big{\lvert}2\Big{\langle}u(x_{i})-v(x_{i})-\big{(}u(x_{j})-v(x_{j})\big{)},u(x_{j})-v(x_{j})\Big{\rangle}+\lVert u(x_{i})-v(x_{i})-(u(x_{j})-v(x_{j}))\rVert^{2}\Big{\rvert}\leq 2C.$
Indeed, after expanding the square, the above is equivalent to
$\big{\lvert}\lVert u(x_{i})-v(x_{i})\rVert^{2}-\lVert u(x_{j})-v(x_{j})\rVert^{2}\big{\rvert}\leq 2C,$
and thus follows by the assumption and the triangle inequality. We see now that for each $j=1,\dotsc,k$
@@
$\Big{\lvert}\Big{\lVert}\sum_{i=1}^{k}t_{i}(u(x_{i})-v(x_{i})-(u(x_{j})-v(x_{j})))\Big{\rVert}^{2}+\sum_{i=1}^{k}t_{i}\lVert u(x_{i})-v(x_{i})-(u(x_{j})-v(x_{j}))\rVert^{2}\Big{\rvert}+3C.$
@@
$\bigg{\lvert}\Big{\lVert}\sum_{i=1}^{k}t_{i}(u(x_{i})-v(x_{i}))\Big{\rVert}^{2}-\delta\bigg{\lvert}\leq 8\big{(}\mathrm{diam}A\big{)}^{2}+3C.$
@@
(4.3) $\bigg{\lvert}\Big{\lVert}\sum_{i=1}^{m}t_{i}(u(x_{i})-v(x_{i}))\Big{\rVert}^{2}-\delta\bigg{\lvert}\leq 8\big{(}\mathrm{diam}A\big{)}^{2}+3C.$
@@
$2\Big{\langle}\Delta_{u}-\Delta_{v},\sum_{i=1}^{m}t_{i}(u(x_{i})-v(x_{i}))\Big{\rangle}+\Big{\lVert}\sum_{i=1}^{m}t_{i}(u(x_{i})-v(x_{i}))\Big{\rVert}^{2}+\lVert\Delta_{u}-\Delta_{v}\rVert^{2}\leq\delta+C.$
Therefore, by (4.3), we get that
$2\Big{\langle}\Delta_{u}-\Delta_{v},\sum_{i=1}^{m}t_{i}(u(x_{i})-v(x_{i}))\Big{\rangle}\leq C+\Big{\lvert}\Big{\lVert}\sum_{i=1}^{m}t_{i}(u(x_{i})-v(x_{i}))\Big{\rVert}^{2}-\delta\Big{\rvert}\leq 8(\mathrm{diam}A)^{2}+4C.$
@@
$-2\lambda\Big{\langle}\Delta_{u}-\Delta_{v},\Delta_{v}\Big{\rangle}+2\Big{\langle}\Delta_{u}-\Delta_{v},\sum_{i,j=1}^{m}t_{i}s_{j}(u(x_{i})-v(x_{i})-u(x_{j})+v(x_{j}))\Big{\rangle}.$
Thus, thanks to the $1$-Lipschitzness of $u$ and of $v$, we get
$-2\lambda\Big{\langle}\Delta_{v},\Delta_{u}-\Delta_{v}\Big{\rangle}\leq 16(\mathrm{diam}A)^{2}+4C.$
Since by (4.5) $\lVert\Delta_{u}\rVert\leq\lVert\Delta_{x}\rVert$, we have
$2\lambda\lVert\Delta_{v}\rVert^{2}\leq 16(\mathrm{diam}A)^{2}+4C+2\lambda\lVert\Delta_{v}\rVert\lVert\Delta_{x}\rVert.$
@@
$\lambda\lVert\Delta_{v}\rVert\leq\frac{8(\mathrm{diam}A)^{2}+2C}{\lVert\Delta_{v}\rVert-\lVert\Delta_{x}\rVert}.$
Similarly as in (4.3) we show that, using (4.2),
$\lvert\lambda^{2}\lVert\Delta_{v}\rVert^{2}-\delta\rvert\leq 8(\mathrm{diam}A)^{2}+3C.$
@@
$\delta\leq\lambda^{2}\lVert\Delta_{v}\rVert^{2}+8(\mathrm{diam}A)^{2}+3C\leq 8(\mathrm{diam}A)^{2}+3C+\bigg{(}\frac{8\,(\mathrm{diam}A)^{2}+2C}{\lVert\Delta_{v}\rVert-\lVert\Delta_{x}\rVert}\bigg{)}^{2}.$
@@ 4.3. Existence of embedding and completion of the proof for $\mathrm{dim}Y\leq 3$
It will follow that (4.6) is also satisfied. Indeed, (4.7) implies trivially that
@@
For any distinct numbers $i,j\in\{1,2,3\}$ we may find $u(x_{i}),u(x_{j})$ in the way described above. We infer by Lemma 4.4 that for any $j\in\{1,2,3\}$ and any non-negative real numbers $(r_{i})_{i\in\{1,2,3\}\setminus\{j\}}$ that sum up to one we have
$\Big{\lVert}v(x_{j})-\sum_{i\in\{1,2,3\}\setminus\{j\}}r_{i}v(x_{i})\Big{\rVert}\leq\Big{\lVert}x_{j}-\sum_{i\in\{1,2,3\}\setminus\{j\}}r_{i}x_{i}\Big{\rVert}.$
@@
Next, we shall construct $u(x_{3})$, so that $u\colon\{x_{1},x_{2},x_{3}\}\to Y$ is an isometry and such that (4.7) holds true. Observe that the possible positions for $u(x_{3})$, for which $u\colon\{x_{1},x_{2},x_{3}\}\to Y$ is an isometry, lie on a circle contained in $Y$. The intersection of the considered circle with $W$ consists of two points $y_{+},y_{-}$. There are two possible positions of $v(x_{3})$ in $W$, given the positions of $v(x_{1}),v(x_{2})$, which we shall denote by $z_{+}$ and $z_{-}$. The situation is illustrated on Figure 1.
@@
Using (4.12), we see that $\lVert v(x_{1})-v(x_{2})\rVert=\lVert x_{1}-x_{2}\rVert$. Thus, in (4.15), we get equalities, which also shows that $u(x_{3})$ satisfying our requirements exists.
@@
Then $v$ satisfies i) of Theorem 1.2, cf. Remark 3.1. We shall specify $C>0$ and $\delta>0$ later. Suppose that we can find an isometry $u\colon A\to Y$ with the required properties. Then
@@
$\lvert\langle u((1,1))-v((1,1)),w\rangle-\delta\rvert=\Big{\lvert}\langle u((1,1))-u((0,0)),w\rangle-\frac{1}{\sqrt{2}}\Big{\rvert}.$
@@
Let us suppose that $X$ is a convex set and $\mathrm{dim}Y\geq 2$. Let $v\colon X\to Y$. Then $v$ is $1$-Lipschitz and affine if and only if for all finite $m\leq\mathrm{dim}Y$, all $x,x_{1},\dotsc,x_{m}\in X$ and all non-negative real numbers $t_{1},\dotsc,t_{m}$ that sum up to one there is
@@
Let us suppose that $X$ is convex and that iii) holds true. By Lemma 4.4 and Lemma 4.8 we see that for $m\leq\min\{2,\mathrm{dim}Y\}$ and all $x,x_{1},\dotsc,x_{m}\in X$ and all non-negative real numbers $t_{1},\dotsc,t_{m}$ that sum up to one we have
@@ 5. Existence of extensions of monotone maps
We need to check that it is a Kirszbraun function, see Definition 2.1. The condition of convexity and finitely lower semi-continuity is clearly satisfied. We need only to check whether the condition (2.1) holds, with $l\leq m$, as $\mathrm{dim}V\leq k-1\leq m-1$. Let $l\leq m$ and $(y_{i},x_{i})_{i=1}^{l}\in V\times X$, $x\in X$ and let $\lambda_{1},\dotsc,\lambda_{l}\geq 0$ be such that $\sum_{i=1}^{l}\lambda_{i}=1$. Then a short calculation readily implies that (2.1) for $\Phi$ is equivalent to
@@
By Theorem 2.3 we infer that for any points $x_{1},\dotsc,x_{k}\in A$, $k\leq m$, and any $x\in X$ the intersection (5.1) of closed, convex sets is non-empty. Indeed, let
$y_{i}=u(x_{i})-v(x_{i})\text{ for }i=1,2,\dotsc,k.$
Then the assumption that $u\colon\{x_{1},\dotsc,x_{k}\}\to Y^{*}$ is monotone is equivalent to
$\Phi(y_{i}-y_{j},x_{i},x_{j})\leq 0\text{ for }i,j=1,2,\dotsc,k.$
Theorem 2.3 implies the existence of $y\in\mathrm{Conv}\{y_{1},\dotsc,y_{k}\}\subset K$ such that
$\Phi(y-y_{i},x_{i},x)\leq 0\text{ for }i=1,2,\dotsc,k,$
which is equivalent to
$\langle y+v(x)-(y_{i}+v(x_{i})),x-x_{i}\rangle\geq 0\text{ for }i=1,2,\dotsc,k.$
@@ 6. Further generalisations
One of the possibilities concerns extensions of maps $u\colon A\to Y$ that satisfy